OpenStudy (anonymous):

int_{}^{}12x^3sqrt{x}

OpenStudy (anonymous):

\[\int\limits_{}^{}12x^3\sqrt{x}\]

OpenStudy (anonymous):

i understand how to integrate each individually but not when multiplying please help.

OpenStudy (anonymous):

write x^3*sqrt(x) as x^3*x^(1/2) = x^(7/2)

OpenStudy (anonymous):

so would it be 108x^(9/2)?

OpenStudy (anonymous):

clearly i dont undrstand

OpenStudy (anonymous):